OpenStudy (anonymous):

Find equation of slope and graph: m=2 (3,5) would the answer be 5=2(3)+b?

OpenStudy (shinalcantara):

point slope form: \[(y-y_1) = m(x-x_1)\]

OpenStudy (shinalcantara):

m is the slope. (x1,y1) is the coordinates. just plug in the given values

OpenStudy (anonymous):

So it would be (y-5)=2(x-3?

OpenStudy (anonymous):

So y-5=2x-6

OpenStudy (shinalcantara):

simplify further by combining the constant terms

OpenStudy (anonymous):

y=2x-1?? (:

OpenStudy (shinalcantara):

yep. if you wanted to have the equation in slope-intercept form, then that's already the answer but if you wanted to have it in standard form then it will be in the form : Ax+By+C=0

OpenStudy (anonymous):

Okay, may I go over another one with you real quick? (:

OpenStudy (shinalcantara):

sure

OpenStudy (anonymous):

Wait first I have to check the ordered pair for the last problem.. How do I check ordered pairs?

OpenStudy (anonymous):

Oh never mind!

OpenStudy (shinalcantara):

say for the question above, all you have to do is plug in the value of x to the equation. check the value of y if it corresponds to the y-coordinate of the given point

OpenStudy (anonymous):

Got it (: So the other one I was going to go over with you is m=2 (-4,-6)

OpenStudy (anonymous):

Soooo.. (y+6)=2(x+4) ???

OpenStudy (anonymous):

Then y+6=2x+8

OpenStudy (shinalcantara):

yep.

OpenStudy (shinalcantara):

go on

OpenStudy (shinalcantara):

combine the constant terms

OpenStudy (anonymous):

So it would be.. y=2x+10!

OpenStudy (shinalcantara):

recheck the value for constant

OpenStudy (anonymous):

Oh wow... Lol oops. y=2x-2

OpenStudy (shinalcantara):

check the sign of the constant xD

OpenStudy (anonymous):

+ oops haha

OpenStudy (shinalcantara):

so the equation would be y = 2x +2

OpenStudy (anonymous):

Yes got it!(: Now I'm going to try one on my own and see if I'm right one second (:

OpenStudy (anonymous):

Okay so for m=2/5 (3,1) I got the equation y=0.4x-.2

OpenStudy (anonymous):

@shinalcantara What is the standard equation of line??

OpenStudy (anonymous):

According to me it is : \(ax + by = c\)

OpenStudy (shinalcantara):

Ax+By+C=0

OpenStudy (anonymous):

No, that is Shinalcantara Equation of line.. :P

OpenStudy (shinalcantara):

hi prrreeeeeeetttttt!!!!! =) :)

OpenStudy (anonymous):

Hi Sheeeenaaaaaaaaaaaaa... :P

OpenStudy (shinalcantara):

:) @waterineyes what are you gonna say to msbambigirl? would you want that equation to be in decimal? O_O

OpenStudy (anonymous):

Which equation??

OpenStudy (anonymous):

I figured it out haha(: I need help with m being undefined and coordinates being (-2,4)!

OpenStudy (shinalcantara):

y=0.4x-.2 that one up up up

OpenStudy (anonymous):

I only said that \(ax + by + c = 0\), you said it is Standard Form but actually it is not..

OpenStudy (shinalcantara):

@msbambygirl "Okay so for m=2/5 (3,1) I got the equation y=0.4x-.2" can you express that not in decimal?

OpenStudy (anonymous):

No o:

OpenStudy (shinalcantara):

O_O what's the standard form of the line then? @waterineyes

OpenStudy (anonymous):

It is : \(ax + by = c\) :) Constant term on Right Side and coefficient of \(x\) should be positive.. :)

OpenStudy (anonymous):

I need help on a different one ;-;

OpenStudy (shinalcantara):

\[(y-y_1) = m(x-x_1)\] \[(y-1) = \frac{ 2 }{ 5 }(x-3)\] \[3(y-1)=2(x-3)\] 3y - 3 = 2x - 6 3y - 2x = -6 + 3 3y - 2x = -3 -----in standard form would this be correct @waterineyes O_O

OpenStudy (shinalcantara):

@msbambygirl for m to be undefined, the change of x must be zero

OpenStudy (anonymous):

So it would be (y-4)=0(x+2)??

OpenStudy (anonymous):

No that is not correct.. :P

OpenStudy (anonymous):

I said, the coefficient of \(x\) should be positive, and not of \(y\).. :P

OpenStudy (anonymous):

\(2x - 3y = 3\), This looks fine.. :)

OpenStudy (shinalcantara):

O_O

OpenStudy (anonymous):

O--O

OpenStudy (shinalcantara):

@msbambygirl for the line to have an undefined slope, it should be a vertical line in that case, the value for your x-coordinate will be constant while that of y is infinity

OpenStudy (anonymous):

So how would I even write that?

OpenStudy (shinalcantara):



OpenStudy (anonymous):

For slope of line as Undefined, y can take any value.. But, the value of \(x\) will remain same for every \(y\)..

OpenStudy (anonymous):

So, the equation of line with having slope undefined is written as : \(x = k\)

OpenStudy (anonymous):

\(k\) is the \(x\) coordinate from where the line touches x-axis.

OpenStudy (anonymous):